# One protein is all you need

### v2

@@ E.3.1 Datasets
For the evaluation of terpene substrate classification, we use the largest available dataset of characterized TPS enzymes from Samusevich et al. (2025) and repurpose the original 5-fold cross-validation schema. We focus on the most challenging TPS sequences, defined as those predicted by the TPS detector, proposed by the dataset authors, with confidence scores below 0.8. This filtering results in 104, 98, 113, 100, 97 examples in the individual folds.
@@
For the test evaluation of subcellular location prediction, we use the setHard dataset constructed by Stärk et al. (2021). The dataset was redundancy-reduced, both within itself and relative to all proteins in DeepLoc (Almagro Armenteros et al. (2017); next paragraph), a standard dataset used for training and validating machine learning models. The setHard dataset contains 490 protein sequences, each annotated with one of ten subcellular location classes, such as “Cytoplasm” or “Nucleus”. Since we use ESM-1b (Rives et al., 2021) in our experiments with the dataset, we further filter the data to 432 sequences that do not exceed a length of 1022 amino acids. This step, consistent with Stärk et al. (2021), ensures that ESM-1b can generate embeddings for all proteins.
@@
For hyperparameter tuning in the subcellular location prediction task, we use the test set from the DeepLoc dataset (Almagro Armenteros et al., 2017). Similar to setHard, DeepLoc assigns labels from one of ten subcellular location classes. The dataset contains 2768 proteins, which we further filter to 2457 sequences that do not exceed a length of 1022 amino acids, ensuring compatibility with the embedding capabilities of ESM-1b. Since setHard was constructed to be independent of DeepLoc, setHard provides a leakage-free source of data for validation.
E.3.2 Metrics
@@
The TPS substrate prediction problem is a 12-class multi-label classification task over possible TPS substrates. Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes. These metrics were used in the original EnzymeExplorer paper (Samusevich et al., 2025). We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
@@
To evaluate the performance of subcellular location prediction methods, we use standard classification metrics as employed in Stärk et al. (2021). Accuracy standardly measures the ratio of correctly classified proteins, while Matthew’s correlation coefficient for multiple classes (MCC) serves as an alternative to the Pearson correlation coefficient for classification tasks (Gorodkin, 2004). The F1-score, the harmonic mean of precision and recall, evaluates performance from a retrieval perspective, balancing the trade-off between false positives and false negatives.
E.3.3 Models
@@
EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (Samusevich et al., 2025). The model consists of two parallel tracks. Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (Meier et al., 2021) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (Jumper et al., 2021). The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
@@
We use Light attention (Stärk et al., 2021) as a representative baseline for subcellular location prediction. Light attention leverages protein embeddings from a language model, which in our case is ESM-1b (Rives et al., 2021). The model processes per-residue embeddings via a softmax-weighted aggregation mechanism, referred to as light attention, which operates with linear complexity relative to sequence length and enables richer aggregation of per-residue information, as opposed to standard mean pooling. We re-train the model using ESM-1b embeddings on the DeepLoc dataset (Almagro Armenteros et al., 2017) using the code from the official GitHub page [^11] https://github.com/HannesStark/protein-localization.
@@
Table A3: Hyperparameters used for adapting ProteinTTT to individual models. The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction. Comma-separated lists show the values used for hyperparameter grid search, while the final values selected for computing the test results are highlighted in bold. Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone. Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A10. Therefore, we used 30 steps in all our experiments. The only exception was ESM3 + ProteinTTT, where the number of steps was set to 50 during initial experiments with different models/tasks conducted in parallel before standardizing the number of steps to 30. Methods marked with an asterisk (“*”) used a slightly different calculation of the loss function. Specifically, the loss was propagated over all tokens, including special and non-masking tokens, while averaging the loss across all tokens simultaneously, rather than first averaging over sequences. This approach was used in the early stages of development, and we provide it in our codebase via loss_kind = ‘‘unnormalized_cross_entropy’’. Please note that MSA Transformer always uses 1 MSA in a batch and the “Batch size” represents the number of sequences in this MSA with the target sequence always present as the first one.
@@
HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
@@
Appendix F Case study details
F.1 Modeling antibody-antigen loops
We download the SAbDab dataset from the official website [^12] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(Dunbar et al., 2014). We apply ProteinTTT to targets with low-confidence ESMFold predictions (pLDDT < 70) and remove sequences longer than 400 residues due to GPU memory limitations. This results in a final set of 175 antibody and 814 antigen chains. We predict the full structures using ESMFold+ProteinTTT (with the same hyperparameters tuned on the CAMEO validation set specified in Table˜A3) and compute LDDT scores against the corresponding PDB structures to assess local errors, which are particularly relevant for loop regions. For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (Chothia & Lesk, 1987) and calculate LDDT on these regions.
F.2 Expanding known structures of viral proteins
We use BFVD version archived/2023_02_v2131313https://bfvd.steineggerlab.workers.dev. This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (Mirdita et al., 2022) with MSAs constructed using Logan (Chikhi et al., 2024), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan. In Figure˜5, we also report pLDDT values for BFVD version archived/2023_02_v1, where structures are simply obtained from ColabFold with MSAs from Logan, i.e., strategy (i). We re-predict structures using ESMFold and ESMFold+ProteinTTT for sequences with length < 450 due to GPU memory constraints. We use the same hyperparameters tuned on the CAMEO validation set, as specified in Table˜A3, with the exception of 20 instead of 30 steps for computational efficiency.
Appendix G Extended results
In this section, we provide additional results on test sets (Section˜G.1), discuss validation performance (Section˜G.2), and analyze the runtime performance of customization (Section˜G.3).
G.1 Detailed test performance
@@
Next, Figure˜A6 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance. This positions ProteinTTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research. While we demonstrate these effects using a protein folding example, we observe a similar distribution of ProteinTTT impact across the tasks.
We also observe that the overall trend of customization with ProteinTTT generally leads to improved performance, with robust consistency across random seeds. However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A8).
G.2 Validation performance
@@
Next, we discuss the hyperparameter optimization. Table˜A3 provides the grid of hyperparameters explored for each model and its size, as well as specifies the optimal hyperparameters suitable for downstream applications. Figure˜A10 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein. While most of reasonable hyperparameter configurations lead to overall improvements when using customization with ProteinTTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting. However, with a reliable predicted confidence measure, such as pLDDT, the appropriate customization step for each protein can be selected to mitigate overfitting. Figure˜A11 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
G.3 Runtime performance
@@
In this section, we demonstrate that customization with ProteinTTT can be done efficiently, with an acceptable computational overhead. Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (Jumper et al., 2021) or AlphaFold3 (Abramson et al., 2024), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
@@
This observation highlights the practical utility of ProteinTTT. For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (Lin et al., 2023). Nevertheless, the original ESMFold has high confidence predictions only for 36% of sequences from the metagenomic database, while the other 392 million sequences remain with low or medium confidence predictions. At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A6 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences). It means that applying ESMFold + ProteinTTT to these remaining sequences could significantly expand the metagenomic atlas characterized by ESMFold. Here, we illustrate this on a similar case study by applying ESMFold + ProteinTTT to more than 300 thousand viral proteins in BFVD (Section˜5.2)
@@
Table A4: ProteinTTT performance on ProteinGym depending on MSA depth. MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜E.2.2). Customization with ProteinTTT primarily improves performance on difficult targets, with low MSA depth. Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
@@
Spearman by MSA depth $\uparrow$
Low depth Medium depth High depth
ESM2 (35M) (Lin et al., 2023) 0.3211 0.2394 0.2707 0.451
ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598
SaProt (35M) (Su et al., 2023) 0.4062 0.3234 0.3921 0.5057
SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091
ESM2 (650M) (Lin et al., 2023) 0.4139 0.3346 0.4063 0.5153
ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075
SaProt (650M) (Su et al., 2023) 0.4569 0.3947 0.4502 0.5448
SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439
ProSST (K=2048) (Li et al., 2024) 0.5068 0.4731 0.5107 0.5749
ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
Table A5: Performance of ProteinTTT on the MaveDB dataset. In this work, we use our newly constructed MaveDB dataset as a validation fold for tuning the ProteinTTT hyper-parameters for fitness prediction. For computational efficiency, we only select a subset of 50 proteins (Section˜E.2.1) and do not run customization across multiple random seeds to estimate standard deviations. The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
Avg. Spearman $\uparrow$
ESM2 (35M) (Lin et al., 2023) 0.4458
@@
Figure A4: Running time of ESMFold + ProteinTTT. For ESMFold and its variants, the median and interquartile ranges of running times on the CAMEO test set are shown using a single NVIDIA A100 GPU. For AlphaFold2, we use estimates from Lin et al. (2023). Specifically, a forward pass through AlphaFold2 is approximately 60 times more computationally expensive than ESMFold (e.g., AlphaFold2 (no MSA; estimate): $2\times 60=120$ seconds), with additional MSA construction taking at least 10 minutes using standard pipelines (AlphaFold2 (estimate): $2\times 60+10\times 60=720$ seconds). ESMFold + ProteinTTT (30 steps) involves LoRA parameter updates, along with forward passes at each customization step to estimate pLDDT and select the structure with the highest predicted confidence. Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜G.2). Overall, ESMFold + ProteinTTT maintains the speed advantage of ESMFold, and is at least an order of magnitude faster than AlphaFold2.
Figure A5: Example of protein fitness prediction upon single-sequence model customization with ProteinTTT. Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (Tsuboyama et al., 2023) (left). ESM2 + ProteinTTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle). The ground-truth fitness data aligns with the customized model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right). Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.

### v3

@@ F.3.1 Datasets
For the evaluation of terpene substrate classification, we use the largest available dataset of characterized TPS enzymes from samusevich2024discovery and repurpose the original 5-fold cross-validation schema. We focus on the most challenging TPS sequences, defined as those predicted by the TPS detector, proposed by the dataset authors, with confidence scores below 0.8. This filtering results in 104, 98, 113, 100, 97 examples in the individual folds.
@@
For the test evaluation of subcellular location prediction, we use the setHard dataset constructed by stark2021light. The dataset was redundancy-reduced, both within itself and relative to all proteins in DeepLoc (almagro2017deeploc; next paragraph), a standard dataset used for training and validating machine learning models. The setHard dataset contains 490 protein sequences, each annotated with one of ten subcellular location classes, such as “Cytoplasm” or “Nucleus”. Since we use ESM-1b (rives2021biological) in our experiments with the dataset, we further filter the data to 432 sequences that do not exceed a length of 1022 amino acids. This step, consistent with stark2021light, ensures that ESM-1b can generate embeddings for all proteins.
@@
For hyperparameter tuning in the subcellular location prediction task, we use the test set from the DeepLoc dataset (almagro2017deeploc). Similar to setHard, DeepLoc assigns labels from one of ten subcellular location classes. The dataset contains 2768 proteins, which we further filter to 2457 sequences that do not exceed a length of 1022 amino acids, ensuring compatibility with the embedding capabilities of ESM-1b. Since setHard was constructed to be independent of DeepLoc, setHard provides a leakage-free source of data for validation.
F.3.2 Metrics
@@
The TPS substrate prediction problem is a 12-class multi-label classification task over possible TPS substrates. Therefore, we assess the quality of the predictions using standard multi-label classification metrics such as mean average precision (mAP) and area under the receiver operating characteristic curve (AUROC) averaged across individual classes. These metrics were used in the original EnzymeExplorer paper (samusevich2024discovery). We report the performance by averaging the metric values concatenated across all validation folds from the 5-fold cross-validation schema.
@@
To evaluate the performance of subcellular location prediction methods, we use standard classification metrics as employed in stark2021light. Accuracy standardly measures the ratio of correctly classified proteins, while Matthew’s correlation coefficient for multiple classes (MCC) serves as an alternative to the Pearson correlation coefficient for classification tasks (gorodkin2004comparing). The F1-score, the harmonic mean of precision and recall, evaluates performance from a retrieval perspective, balancing the trade-off between false positives and false negatives.
F.3.3 Models
@@
EnzymeExplorer is a state-of-the-art method for the classification of terpene synthase (TPS) substrates (samusevich2024discovery). The model consists of two parallel tracks. Given a protein sequence, EnzymeExplorer first computes its ESM-1v embedding (meier2021language) and a vector of similarities to the functional domains of proteins from the training dataset, based on unsupervised domain segmentation of AlphaFold2-predicted structures (jumper2021highly). The ESM-1v embedding and the similarity vector are then concatenated and processed by a separately trained random forest, which predicts TPS substrate class probabilities.
@@
We use Light attention (stark2021light) as a representative baseline for subcellular location prediction. Light attention leverages protein embeddings from a language model, which in our case is ESM-1b (rives2021biological). The model processes per-residue embeddings via a softmax-weighted aggregation mechanism, referred to as light attention, which operates with linear complexity relative to sequence length and enables richer aggregation of per-residue information, as opposed to standard mean pooling. We re-train the model using ESM-1b embeddings on the DeepLoc dataset (almagro2017deeploc) using the code from the official GitHub page [^13] https://github.com/HannesStark/protein-localization.
@@
Table A3: Hyperparameters used for adapting ProteinTTT to individual models. The optimal hyperparameters were estimated using validation datasets corresponding to each of the considered tasks: Fitness prediction, Structure prediction, and Function prediction. Comma-separated lists show the values used for hyperparameter grid search, while the final values selected for computing the test results are highlighted in bold. Low-rank adaptation (LoRA) was only used with ESMFold, containing 3 billion parameters in the ESM2 backbone. Please note that we did not tune the number of customization steps, as adjusting the learning rate and batch size effectively controls the expected performance under the fixed number of steps, as shown in Figure˜A9. Therefore, we used 30 steps in all our experiments. The only exception was ESM3 + ProteinTTT, where the number of steps was set to 50 during initial experiments with different models/tasks conducted in parallel before standardizing the number of steps to 30. Bidirectional methods marked with an asterisk (“*”) used a slightly different calculation of the loss function. Specifically, the loss was propagated over all tokens, including special and non-masking tokens, while averaging the loss across all tokens simultaneously, rather than first averaging over sequences. This approach was used in the early stages of development, and we provide it in our codebase via loss_kind = ‘‘unnormalized_cross_entropy’’. Please note that MSA Transformer always uses 1 MSA in a batch and the “Batch size” represents the number of sequences in this MSA with the target sequence always present as the first one.
@@
ProGen2-small (151M) + ProteinTTT 4e-5, 4e-4, 4e-3 4 4 1, 5, 10, 15, 20 - -
ProGen2-large (2.7B) + ProteinTTT 1e-5 4 4 10, 15, 20 4, 8 8, 16
@@
ESM3 + ProteinTTT 1e-4, 4e-4, 1e-3 2 1, 4, 16 50 (pLDDT) - -
DPLM2 Bit-based + ProteinTTT 4e-6, 4e-5, 4e-4, 4e-3 2, 4, 8 2, 4, 8 10 - -
HelixFold-Single + ProteinTTT 4e-4, 1e-3 4, 8, 16 1 30 (pLDDT) - -
@@
Appendix G Case study details
G.1 Modeling antibody-antigen loops
We download the SAbDab dataset from the official website [^14] https://opig.stats.ox.ac.uk/webapps/sabdab-sabpred/sabdab(dunbar2014sabdab). We apply ProteinTTT to targets with low-confidence ESMFold predictions (pLDDT < 70) and remove sequences longer than 400 residues due to GPU memory limitations. This results in a final set of 175 antibody and 814 antigen chains. We predict the full structures using ESMFold+ProteinTTT (with the same hyperparameters tuned on the CAMEO validation set specified in Table˜A3) and compute LDDT scores against the corresponding PDB structures to assess local errors, which are particularly relevant for loop regions. For antibodies, we evaluate the complete structures, while for complementarity-determining regions (CDRs), we extract the CDR substructures as annotated in SAbDab according to Chothia numbering (chothia1987canonical) and calculate LDDT on these regions.
G.2 Expanding known structures of viral proteins
We use BFVD version archived/2023_02_v2151515https://bfvd.steineggerlab.workers.dev. This version contains maximum-pLDDT structures from predictions generated by two strategies: (i) ColabFold (mirdita2022colabfold) with MSAs constructed using Logan (chikhi2024logan), and (ii) ColabFold with 12 additional recycle steps and MSAs constructed using Logan. In Figure˜5, we also report pLDDT values for BFVD version archived/2023_02_v1, where structures are simply obtained from ColabFold with MSAs from Logan, i.e., strategy (i). We re-predict structures using ESMFold and ESMFold+ProteinTTT for sequences with length < 450 due to GPU memory constraints. We use the same hyperparameters tuned on the CAMEO validation set, as specified in Table˜A3, with the exception of 20 instead of 30 steps for computational efficiency.
Appendix H Extended results
In this section, we provide additional results on test sets (Section˜H.1), discuss validation performance (Section˜H.2), and analyze the runtime performance of customization (Section˜H.3).
H.1 Detailed test performance
@@
Next, Figure˜A7 shows the distribution of ProteinTTT effects: in many cases, customization has minimal impact on performance; often, it leads to substantial improvements; and in rare cases, customization results in a decrease in performance. This positions ProteinTTT as a method for enhancing prediction accuracy, while a comprehensive analysis of its failure modes remains an important direction for future research. While we demonstrate these effects using a protein folding example, we observe a similar distribution of ProteinTTT impact across the tasks.
We also observe that the overall trend of customization with ProteinTTT generally leads to improved performance, with robust consistency across random seeds. However, the progression of the performance curve can be rugged, particularly in classification tasks, where substantial changes in the underlying representations are required to shift the top-predicted class in the discrete probability distribution (Figure˜A12).
H.2 Validation performance
@@
Next, we discuss the hyperparameter optimization. Table˜A3 provides the grid of hyperparameters explored for each model and its size, as well as specifies the optimal hyperparameters suitable for downstream applications. Figure˜A9 demonstrates the trend of hyperparameter tuning with optimal hyperparameter combination balancing underfitting and overfitting to a single target protein. While most of reasonable hyperparameter configurations lead to overall improvements when using customization with ProteinTTT, poorly chosen hyperparameters can have detrimental effects due to rapid overfitting. However, with a reliable predicted confidence measure, such as pLDDT, the appropriate customization step for each protein can be selected to mitigate overfitting. Figure˜A10 demonstrates that when using ESM3 + ProteinTTT with pLDDT-based step selection for protein structure prediction, all hyperparameter configurations result in improved performance compared to the base ESM3 model.
H.3 Runtime performance
@@
In this section, we demonstrate that customization with ProteinTTT can be done efficiently, with an acceptable computational overhead. Specifically, we show that ESMFold, known for being a faster alternative to more performant methods such as AlphaFold2 (jumper2021highly) or AlphaFold3 (abramson2024accurate), still remains in the category of lightweight methods even with ProteinTTT customization (Figure˜A4).
@@
This observation highlights the practical utility of ProteinTTT. For example, ESMFold enabled structural characterization of large metagenomics data (>617 million metagenomic sequences), which would be infeasible with AlphaFold2 (lin2023evolutionary). Nevertheless, the original ESMFold has high confidence predictions only for 36% of sequences from the metagenomic database, while the other 392 million sequences remain with low or medium confidence predictions. At the same time, ESMFold + ProteinTTT enables more accurate predictions compared to the original ESMFold (Figure˜A7 suggests that ESMFold + ProteinTTT significantly improves predictions in almost 40% of challenging sequences). It means that applying ESMFold + ProteinTTT to these remaining sequences could significantly expand the metagenomic atlas characterized by ESMFold. Here, we illustrate this on a similar case study by applying ESMFold + ProteinTTT to more than 300 thousand viral proteins in BFVD (Section˜5.2)
Appendix I Limitations and future work
We see two main limitations of the current version of ProteinTTT, which we discuss in detail below.
Extension to other model types and tasks.
The current form of the method is only applicable to protein language models (PLMs), i.e., Transformer-based (vaswani2017attention) models pre-trained using bidirectional masked language modeling (rives2021biological) or autoregressive next-token prediction (nijkamp2023progen2). Nevertheless, the concept of test-time training can also be extended to many other models in computational biology, which presents exciting opportunities for future research, as our work demonstrates the high potential of this paradigm for the field of computational biology. For instance, our central experiments in Section˜4.1 use ESMFold (lin2023evolutionary), which is known to often underperform (lin2023evolutionary) more specifialized multiple sequence alignment (MSA)-based structure predictors such as AlphaFold2 (jumper2021highly), AlphaFold-Multimer (evans2021protein), AlphaFold3 (abramson2024accurate), or Boltz-2 (passaro2025boltz).
Nevertheless, all of these models can also be extended with test-time training akin to ProteinTTT. AlphaFold2, and subsequently AlphaFold-Multimer, use masked modeling of MSA as one of the training objectives to learn powerful pairwise representations in Evoformer. The Evoformer backbone could therefore be updated at test time to obtain more powerful representation of one input MSA at a time using the ProteinTTT objective (Section˜3.1). While AlphaFold3 and Boltz-2 do not use masked modeling, they can still be customized in a self-supervised way, for example using an optimization through distogram (cho2025boltzdesign1). Implementing the variants of ProteinTTT for the aforementioned models could enable customized structure prediction of protein multimers and protein-ligand complexes.
Beyond structure prediction, test-time customization could also benefit de novo protein design. Our results with autoregressive ProGen2 on fitness prediction suggest that ProteinTTT can improve sequence design (Table˜4). Similarly, although our experiments with ESM3 are currently conducted in the context of structure prediction (Table˜2), ProteinTTT can be straightforwardly applied to ESM3 for protein design tasks such as inverse folding or structure inpainting by applying ProteinTTT to the corresponding ESM3 input tracks. Furthermore, BoltzGen (stark2025boltzgen) can be extended with test-time customization in a manner analogous to Boltz-2, discussed above, due to their shared architecture. Performing ProteinMPNN (dauparas2022robust) customization on part of a protein to guide generation of the remaining structure or its binder, as well as customizing RFdiffusion (watson2023novo) to a target structure for binder design, represent promising opportunities in protein design with the potential for higher success rates.
Better control over failure cases.
The failure modes of ProteinTTT are not yet fully understood. For instance, combining ESMFold with ProteinTTT decreases performance for several proteins in the CAMEO test set (Figure˜A7). A detailed analysis of these cases shows that the degradation can be attributed to ambiguity in the evaluation itself (Figure˜A13). These examples illustrate the challenge of identifying a general reason for the occasional degradation of performance. As discussed in Section˜H.2, confidence functions (such as pLDDT in structure prediction) allow effectively eliminating overfitting to a single protein and thereby mitigating such failure cases, making confidence prediction an essential component of customization.
While confidence functions begin to emerge across tasks, such as fitness prediction (gurev2025variant; nijkamp2023progen2) and inverse folding (shuai2025sidechain), they are not yet universally available for use with ProteinTTT. In particular, for fitness (Section˜4.2) and function (Section˜4.3) prediction, controlling failure cases remains more challenging due to the absence of a reliable confidence metric. This motivates the development of general, task-agnostic, unsupervised confidence measures (for example, perplexity-based estimates (gurev2025variant)) or a dedicated confidence prediction module within ProteinTTT (abramson2024accurate; jumper2021highly). Another promising direction is deriving confidence estimates from mechanistic interpretability of protein language models (hubotter2025specialization; simon2025interplm; zhang2024protein).
@@ Appendix I Limitations and future work
Table A4: ProteinTTT performance on ProteinGym depending on MSA depth. MSA depth reflects the number of available proteins similar to the target protein and, when using large protein language models, can be interpreted as a measure of the representation of similar proteins in the training data (Section˜F.2.2). Customization with ProteinTTT primarily improves performance on difficult targets, with low MSA depth. Standard deviations are calculated over 5 random seeds but are omitted in the right panel for brevity, where the maximum standard deviation does not exceed 0.0004.
Avg. Spearman $\uparrow$ Spearman by MSA depth $\uparrow$ Low depth Medium depth High depth ESM2 (35M) (lin2023evolutionary) 0.3211 0.2394 0.2707 0.451 ESM2 (35M) + ProteinTTT (Ours) 0.3407 ± 0.00014 0.2445 0.3144 0.4598 ProGen2-small (151M) (nijkamp2023progen2) 0.3255 0.2974 0.3136 0.3765 ProGen2-small (151M) + ProteinTTT (Ours) 0.3591 ± 0.0002 0.3319 0.3636 0.3917 ProGen2-large (2.7B) (nijkamp2023progen2) 0.3724 0.3504 0.3808 0.4090 ProGen2-large (2.7B) + ProteinTTT (Ours) 0.3817 ± 0.00158 0.3627 0.3791 0.4152 SaProt (35M) (su2023saprot) 0.4062 0.3234 0.3921 0.5057 SaProt (35M) + ProteinTTT (Ours) 0.4106 ± 0.00004 0.3253 0.3972 0.5091 ESM2 (650M) (lin2023evolutionary) 0.4139 0.3346 0.4063 0.5153 ESM2 (650M) + ProteinTTT (Ours) 0.4153 ± 0.00003 0.3363 0.4126 0.5075 SaProt (650M) (su2023saprot) 0.4569 0.3947 0.4502 0.5448 SaProt (650M) + ProteinTTT (Ours) 0.4583 ± 0.00001 0.3954 0.4501 0.5439 ProSST (K=2048) (li2024prosst) 0.5068 0.4731 0.5107 0.5749 ProSST (K=2048) + ProteinTTT (Ours) 0.5087 ± 0.00004 0.4809 0.5104 0.5750
Table A5: Performance of ProteinTTT on the MaveDB dataset. In this work, we use our newly constructed MaveDB dataset as a validation fold for tuning the ProteinTTT hyper-parameters for fitness prediction. For computational efficiency, we only select a subset of 50 proteins (Section˜F.2.1) and do not run customization across multiple random seeds to estimate standard deviations. The performance shown was calculated by first aggregating correlations per assay, and then per protein (some assays correspond to the same protein).
@@
ESM2 (35M) (lin2023evolutionary) 0.4458
@@
Figure A4: Running time of ESMFold + ProteinTTT. For ESMFold and its variants, the median and interquartile ranges of running times on the CAMEO test set are shown using a single NVIDIA A100 GPU. For AlphaFold2, we use estimates from lin2023evolutionary. Specifically, a forward pass through AlphaFold2 is approximately 60 times more computationally expensive than ESMFold (e.g., AlphaFold2 (no MSA; estimate): $2\times 60=120$ seconds), with additional MSA construction taking at least 10 minutes using standard pipelines (AlphaFold2 (estimate): $2\times 60+10\times 60=720$ seconds). ESMFold + ProteinTTT (30 steps) involves LoRA parameter updates, along with forward passes at each customization step to estimate pLDDT and select the structure with the highest predicted confidence. Disabling pLDDT significantly reduces computational overhead (ESMFold + ProteinTTT (no pLDDT) compared to ESMFold + ProteinTTT), but may require careful parameter tuning (Section˜H.2). Overall, ESMFold + ProteinTTT maintains the speed advantage of ESMFold, and is at least an order of magnitude faster than AlphaFold2.
Figure A6: Example of protein fitness prediction upon single-sequence model customization with ProteinTTT. Fitness predictions from ESM2 (650M) show poor correlation with experimental fitness values in the ProteinGym test set measured by the stability assay “UBR5_HUMAN_Tsuboyama_2023_1I2T” (tsuboyama2023mega) (left). ESM2 + ProteinTTT achieves significantly higher correlation, likely due to improved detection of disruptive mutations in the protein core that impact protein stability (middle). The ground-truth fitness data aligns with the customized model, showing that residues crucial for stability (i.e., having negative mean fitness) are concentrated in the protein core (right). Residue colors represent the mean fitness upon all single-point substitutions (with the exception of several missing mutations in the ground-truth data), with red indicating residues where mutations have detrimental effects on average.
@@
Figure A10: Hyperparameter search for protein structure prediction with ESM3 + ProteinTTT. We conducted a comprehensive grid search based on three key hyperparameters: learning rate (denoted as “lr”), number of gradient accumulation steps (denoted as “grad_steps”; with the batch size of two), and masking strategy (denoted as “mask”). We explored two learning rates, 4e-4 and 1e-3, three gradient accumulation step values of 1, 4, and 16, and five different masking strategies: uniform sampling of 0.05, 0.5, and 1.0 fractions of amino acids, as well as the “beta30” and “betalinear30” distributions proposed in the ESM3 paper (hayes2024simulating). Each row in the table presents the mean TM-score and LDDT metrics with standard deviations across five random seeds on the CAMEO validation fold. The last row, denoted as “No ProteinTTT”, shows the performance of ESM3 without customization. The results indicate that ESM3 + ProteinTTT is robust to the choice of hyperparameters and consistently outperforms the base model across all configurations. We selected the configuration from the last row (excluding “No ProteinTTT”) to compute the results on the test fold. For the hyperparameter search, we used 30 customization steps instead of 50 to reduce computation time.